THE HIDDEN KERNEL OF MOLECULAR QUASI-LINEARITY: QUANTUM MONODROMY

\underline{MANFRED WINNEWISSER}, BRENDA P. WINNEWISSER, IVAN R. MEDVEDEV, FRANK C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210; and STEPHEN C. ROSS, Department of Physics, University of New Brunswick, Fredericton NB E3B5A3.

Chain molecules with one single low-lying bending mode provide a set of model species for the exploration of quantum monodromy in quasi-linear molecules. Recent work on water^{a} and NCNCS^{b} have shown that the topology of the energymomentum maps of such molecules follows closely predictions based on the mathematical concept of non-trivial monodromy. From experimental data and new general semi-rigid bender (GSRB) calculations which extrapolate beyond the existing data for various species, we can now present the topological properties of the bending-rotation energy-momentum maps of a wide range of molecules, from rigidly linear to rigidly bent. It will also be shown that the energy-momentum map for the end-over-end rotational energy, represented by the effective rotational constant B̄, has unexpected properties which, like the bending-rotation energy-momentum map, are robust across the whole set of molecules studied. The molecules discussed are OCCCS, NCCNO, HCNO, OCCCO, ClCNO and BrCNO, NCNCS, HCCNCO, NCNCO and NCSCN.

aN. F. Zobov, S.V. Shirin, O.L. Polyansky, J. Tennyson, P.F. Coheur, P.F. Bernath, M. Carleer, and R. Colin, Chem. Phys.

Lett. 414, 193–197, 2005.

bB.P. Winnewisser, M. Winnewisser, I. R. Medvedev, M. Behnke, F. C. De Lucia, S. C. Ross and J. Koput, Phys. Rev. Lett.

95, 243002, 2005.